OpenStudy (anonymous):

could some one please tell us how to find the answer x. 2^3 * 2^x = 2^2*x We have tried but cant find the way....

OpenStudy (lgbasallote):

NOTE! \[\huge a^b \times a^c = a^{b+c}\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

take 2 common and then solve it

OpenStudy (anonymous):

you do some algebraic operations and you get 2^3+x=2^2x, from the basic algebra we have that if a^b=c^d then where a=c then c=d.So we have x+3=2x, x=3.. you can also solve it with logarithms

OpenStudy (anonymous):

*where a=c then b=d

OpenStudy (anonymous):

get x + 3 = 2x, x=3

OpenStudy (unklerhaukus):

Ah\[2^3 \times 2^x = 2^{2x}\]\[2^{3+x}=2^{2x}\]\[\log_2(2^{3+x})=\log_2(2^{2x})\]\[(3+x)\log_2(2)=2x\log_2(2)\]\[3+x=2x\]\[3=x\]

OpenStudy (anonymous):

what is with the log? \[2^{3+x}=2^{2x}\iff 3+x=2x\iff x=3\]

OpenStudy (anonymous):

the algebraic function a^b=a^c <=> b=c comes from the logarithmic operations decribed above